OpenStudy (kweenlexis):

What is 7.8×105 m/h in kilometers per day? 3.240×1010 km/day 3.240×108 ​ km/day ​ 1.872×106 ​ km/day ​ 1.872×104 km/day ​

OpenStudy (danjs):

Start with the given , and to convert units you can multiply by ratios of 1, notice the units cancel to leave km/day \[\large \frac{ 7.8*10^5~~m }{ 1~~hr } * \frac{ 1~~km }{ 10^{3}~~m }*\frac{ 24~~hr }{ 1~~day } = ...\]

OpenStudy (kweenlexis):

Yeah but how i do this

OpenStudy (danjs):

Multiply the fractions \[\large = \frac{ 7.8*10^5*24 }{ 10^3 }~~\frac{ km }{ day }\]

OpenStudy (kweenlexis):

18720000 for the top 1?

OpenStudy (kweenlexis):

is it C? @DanJS